# Text4Seg: Reimagining Image Segmentation as Text Generation

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have shown exceptional capabilities in vision-language tasks; however, effectively integrating image segmentation into these models remains a significant challenge. In this paper, we introduce Text4Seg, a novel text-as-mask paradigm that casts image segmentation as a text generation problem, eliminating the need for additional decoders and significantly simplifying the segmentation process. Our key innovation is semantic descriptors, a new textual representation of segmentation masks where each image patch is mapped to its corresponding text label. This unified representation allows seamless integration into the auto-regressive training pipeline of MLLMs for easier optimization. We demonstrate that representing an image with semantic descriptors yields competitive segmentation performance. To enhance efficiency, we introduce the Row-wise Run-Length Encoding (R-RLE), which compresses redundant text sequences, reducing the length of semantic descriptors by 74% and accelerating inference by , without compromising performance. Extensive experiments across various vision tasks, such as referring expression segmentation and comprehension, show that Text4Seg achieves state-of-the-art performance on multiple datasets by fine-tuning different MLLM backbones. Our approach provides an efficient, scalable solution for vision-centric tasks within the MLLM framework. Code available at https://github.com/mc-lan/Text4Seg
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) Yin et al. (2023) have successfully extended the capabilities of powerful Large Language Models (LLMs) into the visual domain. Recent advancements demonstrate the remarkable ability of these models to engage in natural language-based human-computer interaction and text-based reasoning over visual inputs Liu et al. (2024c); Lu et al. (2024); Liu et al. (2024a); Bai et al. (2023); Chen et al. (2024). MLLMs have emerged as powerful tools for vision-centric tasks, including image generation Song et al. (2024); Wang et al. (2024b), object detection Wang et al. (2024a); Ma et al. (2024); Zhang et al. (2023) and semantic segmentation Lai et al. (2024); Zhang et al. (2024b). However, seamlessly integrating MLLMs with these tasks, particularly in dense prediction tasks like semantic segmentation, remains challenging due to the intrinsic differences between language and visual modalities.
A straightforward approach adopted by most existing works Lai et al. (2024); Xia et al. (2024); Zhang et al. (2024b); He et al. (2024); Ren et al. (2024); Rasheed et al. (2024); Zhang et al. (2023); Wu et al. (2024) involves appending additional visual decoders (e.g., SAM Kirillov et al. (2023)) to MLLMs, as illustrated in Fig. 1(a). While effective, this combination presents several limitations: 1) it complicates the end-to-end training pipeline with additional loss functions; 2) it requires careful modifications to MLLM architectures, leading to unexpected challenges when scaling up the training. VisionLLM Wang et al. (2024a) attempts to convert segmentation masks into polygon coordinate sequences, as shown in Fig. 1(b). However, the performance is often unsatisfactory, as LLMs may struggle to associate polygon coordinates with shapes, leading to the reintroduction of segmentation-specific decoders in VisionLLMv2 Jiannan et al. (2024). Finding a more effective method to unlock the segmentation capabilities for MLLMs remains crucial. Such method should adhere to the next-token prediction paradigm of MLLMs for easier optimization, require fewer architectural changes for better scalability, and fully leverage text generation capabilities of LLMs.
@@ 2 Related Work
MLLMs are typically developed by enhancing large language models (LLMs) with visual perception modules, which can generate coherent textual conversations grounded in multimodal inputs. For instance, Flamingo Alayrac et al. (2022) introduces the Perceiver Resampler, which connects a pre-trained vision encoder with LLMs for effective few-shot learning. OpenFlamingo Awadalla et al. (2023) and Otter Li et al. (2023a) build upon this architecture with a focus on multi-modal in-context instruction tuning. BLIP-2 Li et al. (2023b) and InstructBLIP Dai et al. (2023) bridge the modality gap using a lightweight Querying Transformer (Q-Former), demonstrating enhanced performance on zero-shot vision-to-language tasks. The LLaVA seires Liu et al. (2024c; a) employs a linear layer or MLP as a modality connector, trained on multimodal language-image instruction-following data generated with GPT-4, showcasing notable capabilities in multimodal chat interactions. They demonstrate impressive capabilities in multimodal chat interactions. In contrast, Qwen-VL Bai et al. (2023) and mPLUG-Owl2 Ye et al. (2024) explore feature compression to a fixed length through cross-attention mechanisms with learnable queries, optimizing computational efficiency. Recent advancements Liu et al. (2024b); Xu et al. (2024); Li et al. (2024a; b; c); Lin et al. (2023) have focused on enhancing visual encoding through high-resolution inputs. For example, LLaVA-UHD Xu et al. (2024) implements an image modularization strategy, segmenting native-resolution images into smaller, variable-sized slices to improve scalability and encoding efficiency. Similarly, LLaVA-NEXT Liu et al. (2024b) and LLaVA-OneVision Li et al. (2024a) utilize the AnyRes scheme to accommodate high-resolution image inputs. In this work, we present Text4Seg to endow existing MLLMs with image segmentation capabilities based on instruction tuning, without necessitating any changes to their architecture.
@@ 3 Methodology
In this section, we begin with an overview of MLLMs in Sec. 3.1. Next, we elaborate on the design of semantic descriptors and row-wise run-length encoding in Sec. 3.2. Finally, we show how to construct visual instruction data to train our proposed Text4Seg in Sec. 3.3.
@@ 3.2 Semantic Descriptors
Our semantic descriptors are inspired by ViT Dosovitskiy et al. (2021), which represents an image as visual tokens. As illustrated in Fig. 3, the process begins by splitting the image into fixed-size patches and flattening them. Each patch is then represented by its corresponding semantic descriptor. A descriptor can be as simple as a semantic label (e.g., “sky,” “sand”), a phrase (e.g., “brown dog”, “black dog”), or even a more complex textual description (e.g., “a dog in the left”) for intricate scenes. This approach encodes an image into a sequence of semantic descriptors of length , which meets the requirements for integrating image segmentation into MLLMs by:
@@
Adopting a text-as-mask paradigm, fully using the text generation capabilities of LLMs for segmentation.
@@ 3.3 Visual Instruction Tuning of Text4Seg
With such pure text response, Text4Seg can be seamlessly integrated with existing MLLMs without any architectural modifications, as shown in Fig. 5. We use Low-Rank Adaptation (LoRA) Hu et al. (2021), to fine-tune the MLLMs on our visual instruction data, using its original auto-regressive training objective . In contrast to existing models Lai et al. (2024); Zhang et al. (2024b); Rasheed et al. (2024), which typically rely on Continued Pre-Training (CPT) with large, mixed datasets to fuse the architectures before fine-tuning on specific downstream tasks, we apply Supervised Fine-Tuning (SFT) directly on the downstream tasks. During inference, to obtain a better pixel-level semantic mask, we optionally apply SAM as the mask refiner with the coarse mask as its prompt.
@@ 4.1 Implementation Details
| Methods | refCOCO | refCOCO+ | refCOCOg | Avg. | |||||
| val | testA | testB | val | testA | testB | val | test | ||
| Specialised Segmentation Models | |||||||||
| LAVT Yang et al. (2022) | 72.7 | 75.8 | 68.8 | 62.1 | 68.4 | 55.1 | 61.2 | 62.1 | 65.8 |
| ReLA Liu et al. (2023a) | 73.8 | 76.5 | 70.2 | 66.0 | 71.0 | 57.7 | 65.0 | 66.0 | 68.3 |
| Generalist Segmentation Models (7B) | |||||||||
| NEXT-Chat Zhang et al. (2023) | 74.7 | 78.9 | 69.5 | 65.1 | 71.9 | 56.7 | 67.0 | 67.0 | 68.9 |
| LISA Lai et al. (2024) | 74.9 | 79.1 | 72.3 | 65.1 | 70.8 | 58.1 | 67.9 | 70.6 | 69.9 |
| PixelLM Ren et al. (2024) | 73.0 | 76.5 | 68.2 | 66.3 | 71.7 | 58.3 | 69.3 | 70.5 | 69.2 |
| AnyRef He et al. (2024) | 76.9 | 79.9 | 74.2 | 70.3 | 73.5 | 61.8 | 70.0 | 70.7 | 72.2 |
| GSVA Xia et al. (2024) | 77.2 | 78.9 | 73.5 | 65.9 | 69.6 | 59.8 | 72.7 | 73.3 | 71.4 |
| LaSagnA Wei et al. (2024) | 76.8 | 78.7 | 73.8 | 66.4 | 70.6 | 60.1 | 70.6 | 71.9 | 71.1 |
| Groundhog Zhang et al. (2024b) | 78.5 | 79.9 | 75.7 | 70.5 | 75.0 | 64.9 | 74.1 | 74.6 | 74.2 |
| GLaMM Rasheed et al. (2024) | 79.5 | 83.2 | 76.9 | 72.6 | 78.7 | 64.6 | 74.2 | 74.9 | 75.6 |
| Text4Seg | 75.0 | 78.6 | 70.1 | 68.4 | 73.4 | 60.0 | 71.5 | 71.7 | 71.1 |
| Text4Seg | 78.8 | 81.5 | 74.9 | 72.5 | 77.4 | 65.9 | 74.3 | 74.4 | 75.0 |
| Text4Seg | 79.3 | 81.9 | 76.2 | 72.1 | 77.6 | 66.1 | 72.1 | 73.9 | 74.9 |
| Text4Seg | 78.0 | 80.9 | 74.6 | 71.6 | 77.3 | 66.0 | 74.8 | 74.7 | 74.7 |
| Text4Seg | 79.2 | 81.7 | 75.6 | 72.8 | 77.9 | 66.5 | 74.0 | 75.3 | 75.4 |
| Generalist Segmentation Models (13B) | |||||||||
| LISA Lai et al. (2024) | 76.0 | 78.8 | 72.9 | 65.0 | 70.2 | 58.1 | 69.5 | 70.5 | 70.1 |
| GSVA Xia et al. (2024) | 78.2 | 80.4 | 74.2 | 67.4 | 71.5 | 60.9 | 74.2 | 75.6 | 72.8 |
| Text4Seg | 80.2 | 82.7 | 77.3 | 73.7 | 78.6 | 67.6 | 74.0 | 75.1 | 76.2 |
| Methods | Validation Set | Test Set A | Test Set B | Avg. | |||
| gIoU | cIoU | gIoU | cIoU | gIoU | cIoU | ||
| Specialised Segmentation Models | |||||||
| LAVT Yang et al. (2022) | 58.4 | 57.6 | 65.9 | 65.3 | 55.8 | 55.0 | 59.7 |
| ReLA Liu et al. (2023a) | 63.6 | 62.4 | 70.0 | 69.3 | 61.0 | 59.9 | 64.4 |
| Generalist Segmentation Models (7B) | |||||||
| LISA Lai et al. (2024) | 61.6 | 61.8 | 66.3 | 68.5 | 58.8 | 60.6 | 62.9 |
| GSVA Xia et al. (2024) | 66.5 | 63.3 | 71.1 | 69.9 | 62.2 | 60.5 | 65.6 |
| Text4Seg | 69.9 | 63.2 | 69.7 | 67.5 | 62.3 | 59.8 | 65.4 |
| Text4Seg | 74.7 | 69.0 | 74.3 | 73.0 | 67.4 | 66.3 | 70.8 |
| Text4Seg | 73.6 | 67.9 | 74.1 | 72.8 | 66.1 | 64.8 | 69.9 |
| Text4Seg | 74.4 | 68.1 | 73.1 | 71.5 | 66.7 | 65.3 | 69.9 |
| Text4Seg | 74.4 | 69.1 | 75.1 | 73.8 | 67.3 | 66.6 | 71.1 |
| Generalist Segmentation Models (13B) | |||||||
| LISA Lai et al. (2024) | 63.5 | 63.0 | 68.2 | 69.7 | 61.8 | 62.2 | 64.7 |
| GSVA Xia et al. (2024) | 68.0 | 64.1 | 71.8 | 70.5 | 63.8 | 61.3 | 66.6 |
| Text4Seg | 74.8 | 69.8 | 75.1 | 74.3 | 68.0 | 67.1 | 71.5 |

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have shown exceptional capabilities in vision-language tasks; however, effectively integrating image segmentation into these models remains a significant challenge. In this paper, we introduce Text4Seg, a novel text-as-mask paradigm that casts image segmentation as a text generation problem, eliminating the need for additional decoders and significantly simplifying the segmentation process. Our key innovation is semantic descriptors, a new textual representation of segmentation masks where each image patch is mapped to its corresponding text label. This unified representation allows seamless integration into the auto-regressive training pipeline of MLLMs for easier optimization. We demonstrate that representing an image with semantic descriptors yields competitive segmentation performance. To enhance efficiency, we introduce the Row-wise Run-Length Encoding (R-RLE), which compresses redundant text sequences, reducing the length of semantic descriptors by 74% and accelerating inference by , without compromising performance. Extensive experiments across various vision tasks, such as referring expression segmentation and comprehension, show that Text4Seg achieves state-of-the-art performance on multiple datasets by fine-tuning different MLLM backbones. Our approach provides an efficient, scalable solution for vision-centric tasks within the MLLM framework. https://github.com/mc-lan/Text4Seg
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) Yin et al. (2023) have successfully extended the capabilities of powerful Large Language Models (LLMs) into the visual domain. Recent advancements demonstrate the remarkable ability of these models to engage in natural language-based human-computer interaction and text-based reasoning over visual inputs Liu et al. (2024c); Lu et al. (2024); Liu et al. (2024a); Bai et al. (2023); Chen et al. (2024). MLLMs have emerged as powerful tools for vision-centric tasks, including image generation Song et al. (2024); Wang et al. (2024d), object detection Wang et al. (2024a); Ma et al. (2024); Zhang et al. (2023) and semantic segmentation Lai et al. (2024); Zhang et al. (2024b); Lan et al. (2024c). However, seamlessly integrating MLLMs with these tasks, particularly in dense prediction tasks like semantic segmentation, remains challenging due to the intrinsic differences between language and visual modalities.
A straightforward approach adopted by most existing works Lai et al. (2024); Xia et al. (2024); Zhang et al. (2024b); He et al. (2024); Ren et al. (2024); Rasheed et al. (2024); Wang et al. (2024c); Zhang et al. (2023); Wu et al. (2024) involves appending additional visual decoders (e.g., SAM Kirillov et al. (2023)) to MLLMs, as illustrated in Fig. 1(a). While effective, this combination presents several limitations: 1) it complicates the end-to-end training pipeline with additional loss functions; 2) it requires careful modifications to MLLM architectures, leading to unexpected challenges when scaling up the training. VisionLLM Wang et al. (2024a) attempts to convert segmentation masks into polygon coordinate sequences, as shown in Fig. 1(b). However, the performance is often unsatisfactory, as LLMs may struggle to associate polygon coordinates with shapes, leading to the reintroduction of segmentation-specific decoders in VisionLLMv2 Jiannan et al. (2024). Finding a more effective method to unlock the segmentation capabilities for MLLMs remains crucial. Such method should adhere to the next-token prediction paradigm of MLLMs for easier optimization, require fewer architectural changes for better scalability, and fully leverage text generation capabilities of LLMs.
@@ 2 Related Work
MLLMs are typically developed by enhancing large language models (LLMs) with visual perception modules, which can generate coherent textual conversations grounded in multimodal inputs. For instance, Flamingo Alayrac et al. (2022) introduces the Perceiver Resampler, which connects a pre-trained vision encoder with LLMs for effective few-shot learning. OpenFlamingo Awadalla et al. (2023) and Otter Li et al. (2023a) build upon this architecture with a focus on multi-modal in-context instruction tuning. BLIP-2 Li et al. (2023b) and InstructBLIP Dai et al. (2023) bridge the modality gap using a lightweight Querying Transformer (Q-Former), demonstrating enhanced performance on zero-shot vision-to-language tasks. The LLaVA seires Liu et al. (2024c; a) employs a linear layer or MLP as a modality connector, trained on multimodal language-image instruction-following data generated with GPT-4, showcasing notable capabilities in multimodal chat interactions. They demonstrate impressive capabilities in multimodal chat interactions. Recent advancements Liu et al. (2024b); Xu et al. (2024); Li et al. (2024a; b; c); Lin et al. (2023) have focused on enhancing visual encoding through high-resolution inputs. For example, LLaVA-UHD Xu et al. (2024) implements an image modularization strategy, segmenting native-resolution images into smaller, variable-sized slices to improve scalability and encoding efficiency. Similarly, LLaVA-NEXT Liu et al. (2024b) and LLaVA-OneVision Li et al. (2024a) utilize the AnyRes scheme to accommodate high-resolution image inputs. In this work, we present Text4Seg to endow existing MLLMs with image segmentation capabilities based on instruction tuning, without necessitating any changes to their architecture.
@@ 3.2 Semantic Descriptors
Our semantic descriptors are inspired by ViT Dosovitskiy et al. (2021), which represents an image as visual tokens. As illustrated in Fig. 3, for simplicity, the example uses visual tokens, the process begins by splitting the image into fixed-size patches and flattening them. Each patch is then represented by its corresponding semantic descriptor. A descriptor can be as simple as a semantic label (e.g., “sky,” “sand”), a phrase (e.g., “brown dog”, “black dog”), or even a more complex textual description (e.g., “a dog in the left”) for intricate scenes. This approach encodes an image into a sequence of semantic descriptors of length , which meets the requirements for integrating image segmentation into MLLMs by:
@@
Adopting a text-as-mask paradigm, using text generation capabilities of LLMs for segmentation.
@@ 3.3 Visual Instruction Tuning of Text4Seg
With such pure text response, Text4Seg can be seamlessly integrated with existing MLLMs without any architectural modifications, as shown in Fig. 5. We use Low-Rank Adaptation (LoRA) Hu et al. (2021), to fine-tune the MLLMs on our visual instruction data, using its original auto-regressive training objective . In contrast to existing models Lai et al. (2024); Zhang et al. (2024b); Rasheed et al. (2024), which typically rely on Continued Pre-Training (CPT) with large, mixed datasets to fuse the architectures before fine-tuning on specific downstream tasks, we apply Supervised Fine-Tuning (SFT) directly on the downstream tasks. During inference, the coarse masks generated by the MLLM demonstrate competitive performance compared to existing methods. To enhance the quality of pixel-level semantic masks, we optionally apply either Conditional Random Fields (CRF) Krähenbühl & Koltun (2011) or the SAM as the mask refiner.
@@ 4.1 Implementation Details
Methods | LLM | refCOCO | refCOCO+ | refCOCOg | Avg. | |||||
val | testA | testB | val | testA | testB | val | test | |||
Specialised Segmentation Models | ||||||||||
ReLA Liu et al. (2023a) | 73.8 | 76.5 | 70.2 | 66.0 | 71.0 | 57.7 | 65.0 | 66.0 | 68.3 | |
HIPIE Wang et al. (2024b) | 78.3 | - | - | 66.2 | - | - | 69.8 | - | - | |
PolyFormer-L Liu et al. (2023b) | 76.0 | 78.3 | 73.3 | 69.3 | 74.6 | 61.9 | 69.2 | 70.2 | 71.6 | |
UNINEXT-L Yan et al. (2023) | 80.3 | 82.6 | 77.8 | 70.0 | 74.9 | 62.6 | 73.4 | 73.7 | 74.4 | |
Generalist Segmentation Models (8B) | ||||||||||
NEXT-Chat Zhang et al. (2023) | Vicuna-7B | 74.7 | 78.9 | 69.5 | 65.1 | 71.9 | 56.7 | 67.0 | 67.0 | 68.9 |
LISA Lai et al. (2024) | Vicuna-7B | 74.9 | 79.1 | 72.3 | 65.1 | 70.8 | 58.1 | 67.9 | 70.6 | 69.9 |
PixelLM Ren et al. (2024) | Vicuna-7B | 73.0 | 76.5 | 68.2 | 66.3 | 71.7 | 58.3 | 69.3 | 70.5 | 69.2 |
AnyRef He et al. (2024) | LLaMA2-7B | 76.9 | 79.9 | 74.2 | 70.3 | 73.5 | 61.8 | 70.0 | 70.7 | 72.2 |
GSVA Xia et al. (2024) | Vicuna-7B | 77.2 | 78.9 | 73.5 | 65.9 | 69.6 | 59.8 | 72.7 | 73.3 | 71.4 |
LaSagnA Wei et al. (2024) | Vicuna-7B | 76.8 | 78.7 | 73.8 | 66.4 | 70.6 | 60.1 | 70.6 | 71.9 | 71.1 |
Groundhog Zhang et al. (2024b) | LLaMA2-7B | 78.5 | 79.9 | 75.7 | 70.5 | 75.0 | 64.9 | 74.1 | 74.6 | 74.2 |
GLaMM Rasheed et al. (2024) | Vicuna-7B | 79.5 | 83.2 | 76.9 | 72.6 | 78.7 | 64.6 | 74.2 | 74.9 | 75.6 |
Text4Seg | DeepSeek-1.3B | 75.0 | 78.6 | 70.1 | 68.4 | 73.4 | 60.0 | 71.5 | 71.7 | 71.1 |
Text4Seg† | DeepSeek-7B | 72.6 | 74.8 | 70.0 | 67.2 | 71.5 | 62.2 | 69.1 | 69.4 | 69.6 |
Text4Seg | DeepSeek-7B | 78.8 | 81.5 | 74.9 | 72.5 | 77.4 | 65.9 | 74.3 | 74.4 | 75.0 |
Text4Seg† | Qwen-7B | 71.3 | 73.7 | 69.6 | 65.9 | 70.4 | 61.9 | 69.3 | 69.3 | 68.9 |
Text4Seg | Qwen-7B | 78.0 | 80.9 | 74.6 | 71.6 | 77.3 | 66.0 | 74.8 | 74.7 | 74.7 |
Text4Seg† | Vicuna-7B | 73.2 | 75.7 | 71.4 | 67.0 | 71.9 | 62.4 | 67.3 | 68.9 | 69.7 |
Text4Seg | Vicuna-7B | 79.3 | 81.9 | 76.2 | 72.1 | 77.6 | 66.1 | 72.1 | 73.9 | 74.9 |
Text4Seg† | InternLM2.5-7B | 73.0 | 75.2 | 70.7 | 67.6 | 72.1 | 62.6 | 68.9 | 70.3 | 70.1 |
Text4Seg‡ | InternLM2.5-7B | 74.7 | 77.4 | 71.6 | 68.5 | 73.6 | 62.9 | 70.7 | 71.6 | 71.4 |
Text4Seg | InternLM2.5-7B | 79.2 | 81.7 | 75.6 | 72.8 | 77.9 | 66.5 | 74.0 | 75.3 | 75.4 |
Generalist Segmentation Models (13B) | ||||||||||
LISA Lai et al. (2024) | Vicuna-13B | 76.0 | 78.8 | 72.9 | 65.0 | 70.2 | 58.1 | 69.5 | 70.5 | 70.1 |
GSVA Xia et al. (2024) | Vicuna-13B | 78.2 | 80.4 | 74.2 | 67.4 | 71.5 | 60.9 | 74.2 | 75.6 | 72.8 |
Text4Seg† | Vicuna-13B | 74.1 | 76.4 | 72.4 | 68.5 | 72.8 | 63.6 | 69.1 | 70.1 | 70.9 |
Text4Seg | Vicuna-13B | 80.2 | 82.7 | 77.3 | 73.7 | 78.6 | 67.6 | 74.0 | 75.1 | 76.2 |
Methods | LLM | Validation Set | Test Set A | Test Set B | Avg. | |||
gIoU | cIoU | gIoU | cIoU | gIoU | cIoU | |||
Specialised Segmentation Models | ||||||||
LAVT Yang et al. (2022) | 58.4 | 57.6 | 65.9 | 65.3 | 55.8 | 55.0 | 59.7 | |
ReLA Liu et al. (2023a) | 63.6 | 62.4 | 70.0 | 69.3 | 61.0 | 59.9 | 64.4 | |
Generalist Segmentation Models (8B) | ||||||||
LISA Lai et al. (2024) | Vicuna-7B | 61.6 | 61.8 | 66.3 | 68.5 | 58.8 | 60.6 | 62.9 |
GSVA Xia et al. (2024) | Vicuna-7B | 66.5 | 63.3 | 71.1 | 69.9 | 62.2 | 60.5 | 65.6 |
Text4Seg | DeepSeek-1.3B | 69.9 | 63.2 | 69.7 | 67.5 | 62.3 | 59.8 | 65.4 |
Text4Seg† | DeepSeek-7B | 70.4 | 65.8 | 68.9 | 69.9 | 63.2 | 63.6 | 67.0 |
Text4Seg | DeepSeek-7B | 74.7 | 69.0 | 74.3 | 73.0 | 67.4 | 66.3 | 70.8 |
Text4Seg† | Qwen-7B | 69.7 | 64.1 | 67.4 | 67.8 | 62.4 | 62.3 | 65.6 |
Text4Seg | Qwen-7B | 74.4 | 68.1 | 73.1 | 71.5 | 66.7 | 65.3 | 69.9 |
Text4Seg† | Vicuna-7B | 69.1 | 64.7 | 69.9 | 70.8 | 62.1 | 62.3 | 66.5 |
Text4Seg | Vicuna-7B | 73.6 | 67.9 | 74.1 | 72.8 | 66.1 | 64.8 | 69.9 |
Text4Seg† | InternLM2.5-7B | 70.0 | 66.1 | 69.4 | 70.9 | 63.1 | 64.1 | 67.3 |
Text4Seg‡ | InternLM2.5-7B | 71.8 | 65.6 | 71.2 | 70.0 | 64.2 | 62.5 | 67.6 |
Text4Seg | InternLM2.5-7B | 74.4 | 69.1 | 75.1 | 73.8 | 67.3 | 66.6 | 71.1 |
Generalist Segmentation Models (13B) | ||||||||
LISA Lai et al. (2024) | Vicuna-13B | 63.5 | 63.0 | 68.2 | 69.7 | 61.8 | 62.2 | 64.7 |
GSVA Xia et al. (2024) | Vicuna-13B | 68.0 | 64.1 | 71.8 | 70.5 | 63.8 | 61.3 | 66.6 |
Text4Seg† | Vicuna-13B | 70.3 | 66.9 | 69.8 | 71.4 | 63.8 | 64.4 | 67.8 |
Text4Seg | Vicuna-13B | 74.8 | 69.8 | 75.1 | 74.3 | 68.0 | 67.1 | 71.5 |